# On Closed-Form Expressions for the Fisher–Rao Distance

### v2

@@ 5 Continuous Distributions
Gumbel $\frac{1}{\sigma}e^{-\frac{x-\mu}{\sigma}}\exp\left(-e^{-\frac{x-\mu}{\sigma}}\right)$ $x\in\mathbb{R}$ $(\mu,\sigma)\in\mathbb{R}\times\mathbb{R}^{*}_{+}$ $\begin{pmatrix}\frac{1}{\sigma^{2}}&\frac{\gamma-1}{\sigma^{2}}\\ \frac{\gamma-1}{\sigma^{2}}&\frac{1}{\sigma^{2}}\text{\tiny$\left((\gamma-1)^{2}+\frac{\pi^{2}}{6}\right)$}\end{pmatrix}$ $\frac{2\pi}{\sqrt{6}}\operatorname{arctanh}\left(\sqrt{\frac{\left[(\mu_{1}-\mu_{2})-(1-\gamma)(\sigma_{1}-\sigma_{2})\right]^{2}+\frac{\pi^{2}}{6}(\sigma_{1}-\sigma_{2})^{2}}{\left[(\mu_{1}-\mu_{2})-(1-\gamma)(\sigma_{1}-\sigma_{2})\right]^{2}+\frac{\pi^{2}}{6}(\sigma_{1}+\sigma_{2})^{2}}}\right)$ [52]
Fréchet $\frac{\lambda}{\beta}\left(\frac{x}{\beta}\right)^{-\lambda-1}e^{-\left(\frac{x}{\beta}\right)^{-\lambda}}$ $x\in\mathbb{R}_{+}^{*}$ $(\beta,\lambda)\in\mathbb{R}_{+}^{*}\times\mathbb{R}_{+}^{*}$ $\begin{pmatrix}\frac{\lambda^{2}}{\beta^{2}}&\frac{1-\gamma}{\beta}\\ \frac{1-\gamma}{\beta}&\frac{1}{\lambda^{2}}\text{\tiny$\left((\gamma-1)^{2}+\frac{\pi^{2}}{6}\right)$}\end{pmatrix}$ $\frac{2\pi}{\sqrt{6}}\operatorname{arctanh}\left(\sqrt{\frac{\left[\log\frac{\beta_{1}}{\beta_{2}}-(1-\gamma)\left(\frac{1}{\lambda_{1}}-\frac{1}{\lambda_{2}}\right)\right]^{2}+\frac{\pi^{2}}{6}\left(\frac{1}{\lambda_{1}}-\frac{1}{\lambda_{2}}\right)^{2}}{\left[\log\frac{\beta_{1}}{\beta_{2}}-(1-\gamma)\left(\frac{1}{\lambda_{1}}-\frac{1}{\lambda_{2}}\right)\right]^{2}+\frac{\pi^{2}}{6}\left(\frac{1}{\lambda_{1}}+\frac{1}{\lambda_{2}}\right)^{2}}}\right)$ [52]
Weibull $\frac{\lambda}{\beta}\left(\frac{x}{\beta}\right)^{\lambda-1}e^{-\left(\frac{x}{\beta}\right)^{\lambda}}$ $x\in\mathbb{R}_{+}^{*}$ $(\beta,\lambda)\in\mathbb{R}_{+}^{*}\times\mathbb{R}_{+}^{*}$ $\begin{pmatrix}\frac{\lambda^{2}}{\beta^{2}}&\frac{\gamma-1}{\beta}\\ \frac{\gamma-1}{\beta}&\frac{1}{\lambda^{2}}\text{\tiny$\left((\gamma-1)^{2}+\frac{\pi^{2}}{6}\right)$}\end{pmatrix}$ $\frac{2\pi}{\sqrt{6}}\operatorname{arctanh}\left(\sqrt{\frac{\left[\log\frac{\beta_{1}}{\beta_{2}}-(1-\gamma)\left(\frac{1}{\lambda_{1}}-\frac{1}{\lambda_{2}}\right)\right]^{2}+\frac{\pi^{2}}{6}\left(\frac{1}{\lambda_{1}}-\frac{1}{\lambda_{2}}\right)^{2}}{\left[\log\frac{\beta_{1}}{\beta_{2}}-(1-\gamma)\left(\frac{1}{\lambda_{1}}-\frac{1}{\lambda_{2}}\right)\right]^{2}+\frac{\pi^{2}}{6}\left(\frac{1}{\lambda_{1}}+\frac{1}{\lambda_{2}}\right)^{2}}}\right)$ [14, 52] [72]
@@
Power function $\theta\alpha^{-\theta}x^{\theta-1}$ $x\in\left]0,\alpha\right]$ $(\theta,\alpha)\in\mathbb{R}_{+}^{*}\times\mathbb{R}_{+}^{*}$ $\begin{pmatrix}\frac{1}{\theta^{2}}&0\\ 0&\frac{\theta^{2}}{\alpha^{2}}\end{pmatrix}$ $2\operatorname{arctanh}\left(\sqrt{\frac{(\theta_{1}\theta_{2}\log(\alpha_{1}/\alpha_{2}))^{2}+(\theta_{1}-\theta_{2})^{2}}{(\theta_{1}\theta_{2}\log(\alpha_{1}/\alpha_{2}))^{2}+(\theta_{1}+\theta_{2})^{2}}}\right)$ particular case in [14]
@@ 5.3 Erlang
An Erlang distribution has p.d.f. $p(x)={\lambda^{k}x^{k-1}e^{-\lambda x}}/{(k-1)!}$, defined for $x\in\mathbb{R}_{+}$, and parametrised by $\lambda\in\mathbb{R}_{+}^{*}$, for a fixed $k\in\mathbb{N}^{*}$. We have $\partial_{\lambda}\ell(\lambda)=\frac{k}{\lambda}-x$, so that the Fisher information is
@@
where we have used that $\mathbb{E}[X]=k/\lambda$ and $\mathbb{E}[X^{2}]=k(k+1)/\lambda^{2}$. The Fisher–Rao distance is then
@@ 5.4.2 Laplace
showing that this metric coincides with the hyperbolic metric 3.1 of the Poincaré half-plane. Using (5.4), we get that the Fisher–Rao distance in this manifold is
@@ 5.4.3 Generalised Gaussian
A generalised Gaussian distribution333 These generalised Gaussian distributions are in a different sense that those considered in [2].  [23, 25, 70], also known as exponential power distribution, is characterised by the p.d.f. $p(x)=\frac{\beta}{2\sigma\Gamma(1/\beta)}\exp\left(-\frac{|x-\mu|^{\beta}}{\sigma}\right)$, defined for $x\in\mathbb{R}$, and parametrised by $(\mu,\sigma)\in\mathbb{R}\times\mathbb{R}_{+}^{*}$, for a fixed $\beta>0$. These can be seen as elliptical distributions with $h(u)=\frac{\beta}{2\Gamma(1/\beta)}\exp(-u^{\beta/2})$, $a_{h}=\frac{\beta\Gamma(2-1/\beta)}{4\Gamma(1+1/\beta)}$, and $b_{h}=\frac{\beta+1}{4}$. Using (5.8), we get the Fisher matrix as
@@ 5.4.6 Student’s $t$
We close this subsection with a remark on the general case of multivariate elliptical distributions. These are distributions of the form $p(\boldsymbol{x})=(\det\Sigma)^{-1/2}h\left((\boldsymbol{x}-\bm{\mu})^{\mathsf{T}}\Sigma^{-1}(\boldsymbol{x}-\bm{\mu})\right)$, for some function $h\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$, defined for $\boldsymbol{x}\in\mathbb{R}^{n}$, and characterised by a vector $\bm{\mu}\in\mathbb{R}^{n}$ and an $n\times n$ positive-definite symmetric matrix $\Sigma\in P_{n}(\mathbb{R})$. Analogously to the univariate case, the set of elliptical distributions for a fixed $h$ form an $\left(n+n(n+1)/2\right)$-dimensional statistical manifold. In the general case, however, no general closed-form expression for the Fisher–Rao distance is known; instead, only expressions for particular cases and bounds for this distance are available [10, 19, 21, 46, 47]. Multivariate Gaussian distributions $p(\boldsymbol{x})=\left((2\pi)^{n}\det\Sigma\right)^{-1/2}\exp\left(-\frac{1}{2}(\boldsymbol{x}-\bm{\mu})^{\mathsf{T}}\Sigma^{-1}(\boldsymbol{x}-\bm{\mu})\right)$ have been particularly studied [5, 14, 17, 18, 36, 56, 66]. Special cases for which the Fisher–Rao distance can be written include fixed mean, fixed covariance, diagonal covariance matrix, and mirrored distributions. The case of diagonal covariance matrix is equivalent to independent components and will be treated in Section 6. Special cases for the multivariate generalised Gaussian distributions have been studied in [70]. The Fisher–Rao distance between zero-mean complex elliptically symmetric distributions has been computed in [12, 13].
@@ 5.6 Inverse Gaussian
An inverse Gaussian distribution  [33, 71] has p.d.f. $p(x)=\sqrt{\frac{\lambda}{2\pi x^{3}}}\exp\left(-\frac{\lambda(x-\mu)^{2}}{2\mu^{2}x}\right)$, defined for $x\in\mathbb{R}_{+}^{*}$, and parametrised by $(\lambda,\mu)\in\mathbb{R}_{+}^{*}\times\mathbb{R}_{+}^{*}$. In this case, we have $\partial_{\lambda}\ell\coloneqq\partial_{\lambda}\ell(\lambda,\mu)=\frac{\lambda(x-\mu)}{\mu^{3}}$ and $\partial_{\mu}\ell\coloneqq\partial_{\mu}\ell(\lambda,\mu)=\frac{1}{2\lambda}-\frac{(x-\mu)^{2}}{2\mu^{2}x}$. The elements of the Fisher matrix are then
@@
$=\frac{1}{4\lambda^{2}}+\frac{1}{\lambda\mu}+\frac{3}{2\mu^{2}}-\left(\frac{1}{2\lambda\mu^{2}}+\frac{1}{\mu^{3}}\right)\mathbb{E}[X]-\left(\frac{1}{2\lambda}+\frac{1}{\mu}\right)\mathbb{E}\left[\frac{1}{X}\right]+\frac{1}{4\mu^{4}}\mathbb{E}\left[X^{2}\right]+\frac{1}{4}\mathbb{E}\left[\frac{1}{X^{2}}\right]$
@@
And, applying Lemma 3.1 yields
@@ 5.7 Extreme-value distributions
Extreme-value distributions [35] are limit distribution for the maxima (or minima) of a sequence of independent and identically distributed random variables. They are usually considered to be of one of three families, all of which can be described in the form of the generalised extreme-value distributions:
@@ 5.7.1 Gumbel
$d_{\mathrm{FR}}\big{(}(\mu_{1},\sigma_{1}),(\mu_{2},\sigma_{2})\big{)}$ $=\frac{\pi}{\sqrt{6}}d_{\mathcal{H}^{2}}\left(\left(\mu_{1}-(1-\gamma)\sigma_{1},\frac{\pi}{\sqrt{6}}\sigma_{1}\right),\left(\mu_{2}-(1-\gamma)\sigma_{2},\frac{\pi}{\sqrt{6}}\sigma_{2}\right)\right)$
@@ 5.7.2 Fréchet
$d_{\mathrm{FR}}\big{(}(\beta_{1},\lambda_{1}),(\beta_{2},\lambda_{2})\big{)}$ $=\frac{\pi}{\sqrt{6}}d_{\mathcal{H}^{2}}\left(\left(\log\beta_{1}-\frac{(1-\gamma)}{\lambda_{1}},\frac{\pi}{\lambda_{1}\sqrt{6}}\right),\left(\log\beta_{2}-\frac{(1-\gamma)}{\lambda_{2}},\frac{\pi}{\lambda_{2}\sqrt{6}}\right)\right)$
@@ 5.7.3 Weibull
$d_{\mathrm{FR}}\big{(}(\beta_{1},\lambda_{1}),(\beta_{2},\lambda_{2})\big{)}$ $=\frac{\pi}{\sqrt{6}}d_{\mathcal{H}^{2}}\left(\left(-\log\beta_{1}-\frac{(1-\gamma)}{\lambda_{1}},\frac{\pi}{\lambda_{1}\sqrt{6}}\right),\left(-\log\beta_{2}-\frac{(1-\gamma)}{\lambda_{2}},\frac{\pi}{\lambda_{2}\sqrt{6}}\right)\right)$
@@ 5.8 Pareto
A Pareto distribution [14, 41] has p.d.f. $p(x)={\theta}\alpha^{\theta}{x^{-(\theta+1)}}$, defined for $x\in\left[\alpha,\infty\right[$ and parametrised by $(\theta,\alpha)\in\mathbb{R}^{*}_{+}\times\mathbb{R}^{*}_{+}$. In this case, the support depends on the parametrisation, thus violating one of the assumptions made in the definition of a statistical manifold (2.1). Nevertheless, it is still possible to compute a Riemannian metric from the Fisher information matrix, as in (2.2)444 As noted in [41], what is not possible is to use the alternative expression (2.3), which would result in a ‘fake metric’. . We thus have $\partial_{\theta}\ell\coloneqq\partial_{\theta}\ell(\theta,\alpha)=\frac{1}{\theta}+\log\alpha-\log x$ and $\partial_{\alpha}\ell\coloneqq\partial_{\alpha}\ell(\theta,\alpha)=\frac{\theta}{\alpha}$. The elements of the Fisher matrix are

### v3

@@ 4 Continuous Distributions
Gumbel $\frac{1}{\sigma}e^{-\frac{x-\mu}{\sigma}}\exp\left(-e^{-\frac{x-\mu}{\sigma}}\right)$ $x\in\mathbb{R}$ $(\mu,\sigma)\in\mathbb{R}\times\mathbb{R}^{*}_{+}$ $\begin{pmatrix}\frac{1}{\sigma^{2}}&\frac{\gamma-1}{\sigma^{2}}\\ \frac{\gamma-1}{\sigma^{2}}&\frac{(\gamma-1)^{2}+\frac{\pi^{2}}{6}}{\sigma^{2}}\end{pmatrix}$ $\frac{2\pi}{\sqrt{6}}\operatorname{arctanh}\left(\sqrt{\frac{\left[(\mu_{1}-\mu_{2})-(1-\gamma)(\sigma_{1}-\sigma_{2})\right]^{2}+\frac{\pi^{2}}{6}(\sigma_{1}-\sigma_{2})^{2}}{\left[(\mu_{1}-\mu_{2})-(1-\gamma)(\sigma_{1}-\sigma_{2})\right]^{2}+\frac{\pi^{2}}{6}(\sigma_{1}+\sigma_{2})^{2}}}\right)$ [18]
Fréchet $\frac{\lambda}{\beta}\left(\frac{x}{\beta}\right)^{-\lambda-1}e^{-\left(\frac{x}{\beta}\right)^{-\lambda}}$ $x\in\mathbb{R}_{+}^{*}$ $(\beta,\lambda)\in\mathbb{R}_{+}^{*}\times\mathbb{R}_{+}^{*}$ $\begin{pmatrix}\frac{\lambda^{2}}{\beta^{2}}&\frac{1-\gamma}{\beta}\\ \frac{1-\gamma}{\beta}&\frac{(\gamma-1)^{2}+\frac{\pi^{2}}{6}}{\lambda^{2}}\end{pmatrix}$ $\frac{2\pi}{\sqrt{6}}\operatorname{arctanh}\left(\sqrt{\frac{\left[\log\frac{\beta_{1}}{\beta_{2}}-(1-\gamma)\left(\frac{1}{\lambda_{1}}-\frac{1}{\lambda_{2}}\right)\right]^{2}+\frac{\pi^{2}}{6}\left(\frac{1}{\lambda_{1}}-\frac{1}{\lambda_{2}}\right)^{2}}{\left[\log\frac{\beta_{1}}{\beta_{2}}-(1-\gamma)\left(\frac{1}{\lambda_{1}}-\frac{1}{\lambda_{2}}\right)\right]^{2}+\frac{\pi^{2}}{6}\left(\frac{1}{\lambda_{1}}+\frac{1}{\lambda_{2}}\right)^{2}}}\right)$ [18]
Weibull $\frac{\lambda}{\beta}\left(\frac{x}{\beta}\right)^{\lambda-1}e^{-\left(\frac{x}{\beta}\right)^{\lambda}}$ $x\in\mathbb{R}_{+}^{*}$ $(\beta,\lambda)\in\mathbb{R}_{+}^{*}\times\mathbb{R}_{+}^{*}$ $\begin{pmatrix}\frac{\lambda^{2}}{\beta^{2}}&\frac{\gamma-1}{\beta}\\ \frac{\gamma-1}{\beta}&\frac{(\gamma-1)^{2}+\frac{\pi^{2}}{6}}{\lambda^{2}}\end{pmatrix}$ $\frac{2\pi}{\sqrt{6}}\operatorname{arctanh}\left(\sqrt{\frac{\left[\log\frac{\beta_{2}}{\beta_{1}}-(1-\gamma)\left(\frac{1}{\lambda_{1}}-\frac{1}{\lambda_{2}}\right)\right]^{2}+\frac{\pi^{2}}{6}\left(\frac{1}{\lambda_{1}}-\frac{1}{\lambda_{2}}\right)^{2}}{\left[\log\frac{\beta_{2}}{\beta_{1}}-(1-\gamma)\left(\frac{1}{\lambda_{1}}-\frac{1}{\lambda_{2}}\right)\right]^{2}+\frac{\pi^{2}}{6}\left(\frac{1}{\lambda_{1}}+\frac{1}{\lambda_{2}}\right)^{2}}}\right)$ [12, 18, 50]
@@
Power function $\theta\beta^{-\theta}x^{\theta-1}$ $x\in\left]0,\beta\right]$ $(\theta,\beta)\in\mathbb{R}_{+}^{*}\times\mathbb{R}_{+}^{*}$ $\begin{pmatrix}\frac{1}{\theta^{2}}&0\\ 0&\frac{\theta^{2}}{\beta^{2}}\end{pmatrix}$ $2\operatorname{arctanh}\left(\sqrt{\frac{(\theta_{1}\theta_{2}\log(\beta_{1}/\beta_{2}))^{2}+(\theta_{1}-\theta_{2})^{2}}{(\theta_{1}\theta_{2}\log(\beta_{1}/\beta_{2}))^{2}+(\theta_{1}+\theta_{2})^{2}}}\right)$ Particular case in [12]
@@
Wishart distribution: $\frac{{(\det X)}^{(n-m-1)/2}\exp\left(-\frac{1}{2}\operatorname{tr}(\Sigma^{-1}X)\right)}{2^{nm/2}(\det\Sigma)^{n/2}\Gamma_{m}(n/2)}$
Inverse Wishart distribution: $\frac{{(\det\Sigma)^{n/2}(\det X)}^{-(n+m+1)/2}\exp\left(-\frac{1}{2}\operatorname{tr}(\Sigma X^{-1})\right)}{2^{nm/2}\Gamma_{m}(n/2)}$
@@ 4.3 Erlang
An Erlang distribution has p.d.f. $p(x)=\frac{\lambda^{k}x^{k-1}e^{-\lambda x}}{(k-1)!}$, defined for $x\in\mathbb{R}_{+}$, and parametrised by $\lambda\in\mathbb{R}_{+}^{*}$, for a fixed $k\in\mathbb{N}^{*}$. We have $\partial_{\lambda}\ell(\lambda)=\frac{k}{\lambda}-x$, so that the Fisher information is
@@
where we have used that $\mathbb{E}[X]=\frac{k}{\lambda}$ and $\mathbb{E}[X^{2}]=\frac{k(k+1)}{\lambda^{2}}$. The Fisher–Rao distance is then
@@ 4.4.2 Laplace
showing that this metric coincides with the hyperbolic metric (2.1) of the Poincaré half-plane. Using (4.4), we get that the Fisher–Rao distance in this manifold is
@@ 4.4.3 Generalised Gaussian
A generalised Gaussian distribution333 These generalised Gaussian distributions are in a different sense that those considered in [52].  [53, 51, 22], also known as exponential power distribution, is characterised by the p.d.f. $p(x)=\frac{\beta}{2\sigma\Gamma(1/\beta)}\exp\left(-\frac{|x-\mu|^{\beta}}{\sigma}\right)$, defined for $x\in\mathbb{R}$, and parametrised by $(\mu,\sigma)\in\mathbb{R}\times\mathbb{R}_{+}^{*}$, for a fixed $\beta>0$. These can be seen as elliptical distributions with $h(u)=\frac{\beta}{2\Gamma(1/\beta)}\exp(-u^{\beta/2})$, $a_{h}=\beta\frac{\Gamma(2-1/\beta)}{\Gamma(1+1/\beta)}$, and $b_{h}=\beta$. Using (4.8), we get the Fisher matrix as
@@ 4.4.6 Student’s $t$
We close this subsection with a remark on the general case of multivariate elliptical distributions. These are distributions of the form $p(\boldsymbol{x})=(\det\Sigma)^{-1/2}h\left((\boldsymbol{x}-\bm{\mu})^{\mathsf{T}}\Sigma^{-1}(\boldsymbol{x}-\bm{\mu})\right)$, for some function $h\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$, defined for $\boldsymbol{x}\in\mathbb{R}^{n}$, and characterised by a vector $\bm{\mu}\in\mathbb{R}^{n}$ and an $n\times n$ positive-definite symmetric matrix $\Sigma\in P_{n}(\mathbb{R})$. Analogously to the univariate case, the set of elliptical distributions for a fixed $h$ forms an $\left(n+\frac{n(n+1)}{2}\right)$-dimensional statistical manifold. In the general case, however, no general closed-form expression for the Fisher–Rao distance is known; instead, only expressions for particular cases and bounds for this distance are available [54, 55, 56, 17, 57]. Multivariate Gaussian distributions $p(\boldsymbol{x})=\left((2\pi)^{n}\det\Sigma\right)^{-1/2}\exp\left(-\frac{1}{2}(\boldsymbol{x}-\bm{\mu})^{\mathsf{T}}\Sigma^{-1}(\boldsymbol{x}-\bm{\mu})\right)$ have been particularly studied [10, 12, 58, 13, 59, 20, 60]. Special cases for which the Fisher–Rao distance can be written include: fixed mean, fixed covariance, diagonal covariance matrix, and mirrored distributions. The case of diagonal covariance matrix is equivalent to independent components and will be treated in Section 5. Special cases for the multivariate generalised Gaussian distributions have been studied in [22]. The Fisher–Rao distance between zero-mean complex elliptically symmetric distributions has been computed in [33, 61].
@@ 4.6 Inverse Gaussian
An inverse Gaussian distribution  [48, 23] has p.d.f. $p(x)=\sqrt{\frac{\lambda}{2\pi x^{3}}}\exp\left(-\frac{\lambda(x-\mu)^{2}}{2\mu^{2}x}\right)$, defined for $x\in\mathbb{R}_{+}^{*}$, and parametrised by $(\lambda,\mu)\in\mathbb{R}_{+}^{*}\times\mathbb{R}_{+}^{*}$. In this case, we have $\partial_{\lambda}\ell\coloneqq\partial_{\lambda}\ell(\lambda,\mu)=\frac{1}{2\lambda}-\frac{(x-\mu)^{2}}{2\mu^{2}x}$ and $\partial_{\mu}\ell\coloneqq\partial_{\mu}\ell(\lambda,\mu)=\frac{\lambda(x-\mu)}{\mu^{3}}$. The elements of the Fisher matrix are then
@@
$=\frac{1}{4\lambda^{2}}+\frac{1}{\lambda\mu}+\frac{3}{2\mu^{2}}-\left(\frac{1}{2\lambda\mu^{2}}+\frac{1}{\mu^{3}}\right)\mathbb{E}[X]$
¨$-\left(\frac{1}{2\lambda}+\frac{1}{\mu}\right)\mathbb{E}\left[\frac{1}{X}\right]+\frac{1}{4\mu^{4}}\mathbb{E}\left[X^{2}\right]+\frac{1}{4}\mathbb{E}\left[\frac{1}{X^{2}}\right]$
@@
Applying Lemma 2.1 yields
@@ 4.7 Extreme-value distributions
Extreme-value distributions [62] are limit distribution for the maxima (or minima) of a sequence of i.i.d. random variables. They are usually considered to be of one of three families, all of which can be described in the form of the generalised extreme-value distributions:
@@ 4.7.1 Gumbel
$d_{\mathrm{FR}}\big{(}(\mu_{1},\sigma_{1}),(\mu_{2},\sigma_{2})\big{)}$
$=\frac{\pi}{\sqrt{6}}d_{\mathcal{H}^{2}}\left(\left(\mu_{1}-(1-\gamma)\sigma_{1},\frac{\pi}{\sqrt{6}}\sigma_{1}\right),\left(\mu_{2}-(1-\gamma)\sigma_{2},\frac{\pi}{\sqrt{6}}\sigma_{2}\right)\right)$
@@ 4.7.2 Fréchet
$d_{\mathrm{FR}}\big{(}(\beta_{1},\lambda_{1}),(\beta_{2},\lambda_{2})\big{)}$
$=\frac{\pi}{\sqrt{6}}d_{\mathcal{H}^{2}}\left(\left(\log\beta_{1}-\frac{(1-\gamma)}{\lambda_{1}},\frac{\pi}{\lambda_{1}\sqrt{6}}\right),\left(\log\beta_{2}-\frac{(1-\gamma)}{\lambda_{2}},\frac{\pi}{\lambda_{2}\sqrt{6}}\right)\right)$
@@ 4.7.3 Weibull
$d_{\mathrm{FR}}\big{(}(\beta_{1},\lambda_{1}),(\beta_{2},\lambda_{2})\big{)}$
$=\frac{\pi}{\sqrt{6}}d_{\mathcal{H}^{2}}\left(\left(-\log\beta_{1}-\frac{(1-\gamma)}{\lambda_{1}},\frac{\pi}{\lambda_{1}\sqrt{6}}\right),\left(-\log\beta_{2}-\frac{(1-\gamma)}{\lambda_{2}},\frac{\pi}{\lambda_{2}\sqrt{6}}\right)\right)$
@@ 4.8 Pareto
A Pareto distribution [12, 14] has p.d.f. $p(x)={\theta}\alpha^{\theta}{x^{-(\theta+1)}}$, defined for $x\in\left[\alpha,\infty\right[$ and parametrised by $(\theta,\alpha)\in\mathbb{R}^{*}_{+}\times\mathbb{R}^{*}_{+}$. In this case, the support depends on the parametrisation, thus violating one of the assumptions made in the definition of a statistical manifold (1.1). Nevertheless, it is still possible444 As noted in [14], what is not possible is to use the alternative expression (1.3), which would result in a ‘fake metric’. to compute a Riemannian metric from the Fisher information matrix, as in (1.2). We thus have $\partial_{\theta}\ell\coloneqq\partial_{\theta}\ell(\theta,\alpha)=\frac{1}{\theta}+\log\alpha-\log x$ and $\partial_{\alpha}\ell\coloneqq\partial_{\alpha}\ell(\theta,\alpha)=\frac{\theta}{\alpha}$. The elements of the Fisher matrix are